HUBBLE SPACE TELESCOPE DAILY REPORT #1145 PERIOD COVERED: 0700 EST Daily Status Report as of 144/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/BL 5649 (SMOV FOS/FGS Fine Alignment Paired Apertures, Repeat) The Faint Object Spectrograph (Blue Side) used STAR-07362 to make precise measurements of the sizes and locations of the FOS apertures. Many of the observations were received and displayed in real-time. The target was reported to be well-centered. Some of the data remains of the Science Tape Recorder, and will be analyzed later. No anomalies were noted. 1.2 Completed WF/PC-2 5454 (Ellipticals with Kinematically Distinct Cores: Cycle 4 Medium) The WF/PC-2 was used to make observations of the small elliptical galaxy NGC-1427 which contains a massive core which is kinematically distinct from the rest of the galaxy. The observations were executed as scheduled. The baseline guide star acquisition failed (See Operations Note 2.1), and the acquisition failed over to single course track with FGS #1. The search limit was exceeded for FGS #1. 1.3 Completed WF/PC-2 5475 (A Snapshot Survey for Blue Stragglers in Magellanic Cluster Cores) The WF/PC-2 was used to make short exposure snapshot images of the blue straggler stars located in populous cluster NGC-1415. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 5561 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats) The WF/PC-2 was used to make a internal dark calibration image to monitor flat field stability and the stability of the analog-to-digital converters. The observation was executed as scheduled, and no anomalies were noted. 1.5 Completed S/C 4974 (DF-224 FS Installation 9.0 and Higher) The installation of Flight Software 9.3 was executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 3 FGS Acquisition failure at 144/0504Z FGS #2 failed to acquire, and FGS #1 failed to Course Track. Scheduled Reacquisitions : 2 Successful : 2 2.2 FHST Updates: Scheduled :12 Successful :12 2.3 Significant HSTARs: None 2.4 Operations Notes: No significant operational problems were encountered during the past twenty-four hours. DF-224 Flight Software Version 9.3 was uplinked to the spacecraft successfully. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete S/C 4974 (DF-224 FS Installation 9.0 and Higher) 3.2 Complete WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) 3.3 Complete FOS/BL 5418 (The Origin of the Ultraviolet Spectral Components) 3.4 Complete WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) 3.4 Complete WF/PC-2 5369 (HST Medium-Deep Survey: Cycle 4 Medium, High Galactic Latitude) 3.5 Complete WF/PC-2 5561 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 2: Flats) 3.6 Complete FOS/BL 5539 (Cycle 4 FOS Photometric Monitor) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-